# Assimilative Causal Inference

### v1

@@ C.2 Principle of Nil Conditional Assimilative Causality for CGNSs
$\overleftarrow{\mathrm{d}(\boldsymbol{\mu}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}-\boldsymbol{\mu}_{\text{{f}}|\mathbf{x}_{\text{{B}}}})}(t)=-(\boldsymbol{\Lambda}^{\mathbf{y}}+\mathbf{B}_{|\mathbf{x}_{\text{{B}}}}\mathbf{R}_{\text{{f}}|\mathbf{x}_{\text{{B}}}}^{-1})(\boldsymbol{\mu}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}-\boldsymbol{\mu}_{\text{{f}}|\mathbf{x}_{\text{{B}}}})\mathrm{d}t,\quad T\geq t\geq 0,$ (3.7)
@@
$\boldsymbol{\mu}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}(t)=\boldsymbol{\mu}_{\text{{f}}|\mathbf{x}_{\text{{B}}}}(t),\quad t\in[0,T].$
@@
$\mathbf{R}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}(t)=\mathbf{R}_{\text{{f}}|\mathbf{x}_{\text{{B}}}}(t),\quad t\in[0,T],$
@@
$p_{t}^{\text{{s}}|\mathbf{x}_{\text{{B}}}}(\mathbf{y}|\mathbf{x}_{\text{{A}}})=p_{t}^{\text{{f}}|\mathbf{x}_{\text{{B}}}}(\mathbf{y}|\mathbf{x}_{\text{{A}}}),\quad t\in[0,T],$
[…]
@@
When $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{A}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})\equiv\mathbf{0}_{k_{\text{A}}\times k_{\text{B}}}$, the interpretation of “assigning infinite uncertainty to $\mathbf{x}_{\text{B}}$’s marginal likelihood” (Section AA.4) becomes straightforward. In this case, $\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty$ simply corresponds to $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})^{-1}\to\mathbf{0}_{k_{\text{B}}\times k_{\text{B}}}$ for all $t$ and $\mathbf{x}$, since $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})$ becomes block-diagonal. For general turbulent systems or when this condition fails, the interpretation of this limiting procedure requires more care. Here, $\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty$ demands rigorous analysis of the structure of $\mathbf{x}_{\text{B}}$’s marginal likelihood and its impact on the Kalman gain operators. Only through such analysis can we properly nullify $\mathbf{x}_{\text{B}}$’s influence when testing for conditional assimilative causal links $\big{(}\mathbf{y}(t)\rightarrow\mathbf{x}_{\text{A}}\big{)}\big{|}\mathbf{x}_{\text{B}}$. We defer this detailed investigation to future work.
@@
When $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{A}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})\equiv\mathbf{0}_{k_{\text{A}}\times k_{\text{B}}}$ in a CGNS, implementing the condition within the generalized ACI framework $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})^{-1}\to\mathbf{0}_{k_{\text{B}}\times k_{\text{B}}}$ (Section AA.4) leads to an equivalent reduced system. In this reduced system, the state estimation of $\mathbf{y}(t)$ becomes governed by the CGNS defined in (3.2a) and (3.2c), where $(\mathbf{x}_{\text{A}},\mathbf{y})$ form the state variables, with $\mathbf{x}_{\text{B}}$ being reduced to a deterministic forcing term defined by its observed values.
@@
This section includes three numerical studies that exploit the ACI framework to study complex dynamical systems with intermittency, non-Gaussian features, regime switching, and extreme events. All test cases employ CGNS models, leveraging their analytical tractability while capturing these nonlinear phenomena.
D.1 A Nonlinear Dyad Model with Extreme Events
@@
For completeness, we restate the nonlinear dyad model described in the main text:
$\frac{{\mathrm{d}}y}{{\mathrm{d}}t}$ $=-d_{y}y-\gamma x^{2}+f_{y}+\sigma_{y}\dot{W}_{y}.$ (4.1b)
@@
This is a reduced-order conceptual model for atmospheric variability. It has been used to analyze the effects of various coarse-grained procedures on processes exhibiting intermittency, large-scale bifurcations, and microscale phase transitions. It is defined by an energy-conserving condition on its quadratic nonlinearities [16]. The coupling parameter $\gamma>0$ plays a crucial role by ensuring significant positive $y$ values ($y>d_{x}/\gamma$) trigger extreme events in $x$.
@@
Figure 4.1 presents the data assimilation results for the dyad model in (4.1). The key distinction between the filter and smoother distributions in estimating $y$ occurs prior to extreme events in $x$. This behavior is expected since the filter, operating without knowledge of future observations, cannot fully anticipate the triggering mechanism in $y$. In contrast, the smoother benefits from future extreme event data, enabling more accurate state estimation of y with reduced uncertainty.
@@
Figure 4.1: Data assimilation of the dyad model (4.1). Panel (a): A single realization of the observed variable $x$. Panel (b): The true hidden signal $y$ (blue) alongside the posterior mean estimates from filtering (green) and smoothing (red), where the smoother is the complete smoother using all the information in future. The dashed line marks the anti-damping threshold, above which the net damping $-d_{x}+\gamma y$ in (4.1a) becomes positive. Panel (c): Posterior variance of the filtered and smoothed estimates of $y$.
@@
Figure 4.2 displays the ACI and CIR analyses. Several important patterns emerge:
@@
  • First, the ACI value reaches its maximum during $y$’s strongest anti-damping phase, corresponding to $y$’s peak instantaneous influence on $x$.
@@
  • Second, the subjective CIR (shown by deep red shading in Panel (b)) attains its highest values slightly before $y>d_{x}/\gamma$ (initiating an extreme event) and for small $\epsilon$ thresholds (indicating a longer-range influence). This reveals that extreme events develop gradually, with triggering conditions established well in advance. Notably, the objective CIR’s temporal extent does not reach the actual peak of extreme events. This finding mirrors the decorrelation time (the integration of the autocorrelation function) in complex dynamical systems, where the true causal influence often persists weakly beyond the formal timescale indicated by the objective function.
@@
  • Third, and most significantly, the objective CIR shortens as the system approaches $y$’s positive peak (corresponding to $x$’s extreme event buildup phase). At this stage, the filter can reliably detect the emerging pattern without requiring future information. This transition naturally partitions the time series at each extreme event, marking distinct dynamical regimes: a build-up phase with long-range dependence (where future information improves estimation) and an event phase where the trajectory becomes locally predictable. The short CIR also persists during $y$’s demise, where $x$ is the driving factor behind the system dynamics (high signal-to-noise ratio) and controls $y$ via $-\gamma x^{2}$.
@@
Figure 4.2: ACI values and CIRs for the nonlinear dyad model (4.1) from $y$ to $x$ as functions of time. Panel (a): Time series of $x$ (magenta) and $y$ (blue), with the objective CIR depicted as whiskers extending forward in time from each $y(t)$. The dashed horizontal line marks the anti-damping threshold $d_{x}/\gamma$. Panel (b): Subjective CIR (shaded region, logarithmically scaled) as a function of the threshold $\epsilon$ (logarithmic, reversed $y$-axis). Panel (c): ACI values from $y$ to $x$ over time.
D.2 A Noisy Predator-Prey Model
The predator-prey model (also known as the Lotka-Volterra model) is fundamentally important across scientific disciplines as it captures the universal dynamics of interacting populations through simple yet powerful mathematics [6]. Originating in ecology to explain cyclical fluctuations between species like lynx and hares, its core principles have been adapted to model diverse systems, from disease spread in epidemiology to competition in economics and even chemical oscillations. The nonlinear feedback mechanisms in the model provide crucial insights into understanding stability, resilience, and emergent patterns of many natural and scientific problems.
Let us consider a noisy version of the predator-prey model:
$\frac{{\mathrm{d}}x}{{\mathrm{d}}t}$ $=\beta xy-\alpha x+\sigma_{x}\dot{W}_{x},$ (4.2a)
$\frac{{\mathrm{d}}y}{{\mathrm{d}}t}$ $=\gamma y-\delta xy+\sigma_{y}\dot{W}_{y},$ (4.2b)
@@ D.2 A Noisy Predator-Prey Model
Figure 4.3 displays the data assimilation results. Panels (a)–(c) and (d)–(f) show the results by observing $y$ (recovering $x$) and observing $x$ (recovering $y$), respectively. The state estimation is more informative when the observed signal has a large value, corresponding to when the estimated variable induces an extreme event in the former’s evolution. Additionally, the uncertainty reduction in the smoother related to the filter is more significant when the predator $x$ is the observed variable, which accounts for the choice of the coupling, quadratic feedback parameter values; $\delta>\beta$.
Figure 4.3: Data assimilation of the noisy predator-prey model (4.2). Panels (a)–(c) and (d)–(f) show the results by observing $y$ (recovering $x$) and observing $x$ (recovering $y$), respectively. The dashed lines in Panels (b) and (e) indicate the anti-damping threshold values.
Figure 4.4 presents the ACI value (instantaneous causal strength) and CIR (influence duration) between predator ($x$) and prey ($y$) populations. The interaction exhibits two distinct regimes:
@@
Figure 4.4: ACI values and CIRs for the noisy predator-prey model (4.2). Only the objective CIRs are shown. Panels (a)–(b) and (c)–(d) show the results from $x(t)$ to $y$ and from $y(t)$ to $x$, respectively. The dashed lines in Panels (a) and (b) indicate the anti-damping threshold values in the equations of $y$ and $x$, respectively.
D.3 A Stochastic Model Capturing the El Niño-Southern Oscillation (ENSO) Diversity
@@
As was described in the main text, although few models can accurately capture ENSO diversity, a recently developed stochastic conceptual model successfully reproduces its diverse behaviors and non-Gaussian statistics [4]. This model has been highlighted in a recent review [22], making it a suitable testbed for studying El Niño diversity. The model consists of six state variables: ocean zonal current in the CP ($u$), western Pacific (WP) thermocline depth ($h_{W}$), CP SST ($T_{C}$), EP SST ($T_{E}$), atmospheric winds ($\tau$, intraseasonal), and decadal variation ($I$, decadal). The variables ($u$, $h_{W}$, $T_{C}$, and $T_{E}$) operate on interannual timescales and model the anomalies away from the corresponding climatology. As a nonlinear system with state-dependent noise, the model generates extreme events and intermittency. Note that the decadal variable represents the strength of the Walker circulation. The two SST variables ($T_{C}$ and $T_{E}$) allow reconstruction of spatiotemporal SST patterns across the equatorial Pacific, providing an intuitive way to identify different ENSO event types. The model reads:
$\frac{{\mathrm{d}}u}{{\mathrm{d}}t}$ $=-ru-\delta_{u}\frac{T_{C}+T_{E}}{2}+\beta_{u}(I)\tau+\sigma_{u}\dot{W}_{u},$ (4.3a)
$\frac{{\mathrm{d}}h_{W}}{{\mathrm{d}}t}$ $=-rh_{W}-\delta_{h}\frac{T_{C}+T_{E}}{2}+\beta_{h}(I)\tau+\sigma_{h}\dot{W}_{h},$ (4.3b)
$\frac{{\mathrm{d}}T_{C}}{{\mathrm{d}}t}$ $=\big{(}r_{C}-c_{1}(t,T_{C})\big{)}T_{C}+\zeta_{C}T_{E}+\gamma_{C}h_{W}+\sigma(I)u+C_{u}+\beta_{C}(I)\tau+\sigma_{C}\dot{W}_{C},$ (4.3c)
$\frac{{\mathrm{d}}T_{E}}{{\mathrm{d}}t}$ $=\big{(}r_{E}-c_{2}(t)\big{)}T_{E}-\zeta_{E}T_{C}+\gamma_{E}h_{W}+\beta_{E}(I)\tau+\sigma_{E}\dot{W}_{E},$ (4.3d)
$\frac{{\mathrm{d}}\tau}{{\mathrm{d}}t}$ $=-d_{\tau}\tau+\sigma_{\tau}(t,T_{C})\dot{W}_{\tau},$ (4.3e)
$\frac{{\mathrm{d}}I}{{\mathrm{d}}t}$ $=-\lambda(I-m)+\sigma_{I}(I)\dot{W}_{I}.$ (4.3f)
@@
While the full model details are available in [4], its basic mechanism can be summarized as follows. EP El Niño events are typically associated with a strong thermocline buildup in the WP, whereas CP EN and La Niña events are primarily triggered by advective processes. In the latter case, when $\sigma(I)\propto I$ is larger, indicating a strengthening Walker circulation, the advection term $u$ becomes dynamically significant in driving $T_{C}$. Unlike the standard discharge-recharge oscillator model, which couples $h_{W}$ and $T_{E}$, the $T_{C}$ variable acts as a transitional component that is expected to have a more direct influence on $T_{E}$.

### v2

@@ C.2 Principle of Nil Conditional Assimilative Causality for CGNSs
$\overleftarrow{\mathrm{d}(\boldsymbol{\mu}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}-\boldsymbol{\mu}_{\text{{f}}|\mathbf{x}_{\text{{B}}}})}(t)=-(\boldsymbol{\Lambda}^{\mathbf{y}}+\mathbf{B}_{|\mathbf{x}_{\text{{B}}}}\mathbf{R}_{\text{{f}}|\mathbf{x}_{\text{{B}}}}^{-1})(\boldsymbol{\mu}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}-\boldsymbol{\mu}_{\text{{f}}|\mathbf{x}_{\text{{B}}}})\mathrm{d}t,\qquad T\geq t\geq 0,$ (3.7)
@@
$\boldsymbol{\mu}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}(t)=\boldsymbol{\mu}_{\text{{f}}|\mathbf{x}_{\text{{B}}}}(t),\qquad t\in[0,T].$
@@
$\mathbf{R}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}(t)=\mathbf{R}_{\text{{f}}|\mathbf{x}_{\text{{B}}}}(t),\qquad t\in[0,T],$
@@
$p_{t}^{\text{{s}}|\mathbf{x}_{\text{{B}}}}(\mathbf{y}|\mathbf{x}_{\text{{A}}})=p_{t}^{\text{{f}}|\mathbf{x}_{\text{{B}}}}(\mathbf{y}|\mathbf{x}_{\text{{A}}}),\qquad t\in[0,T],$
which in turn, by (1.14), establishes $\big(\mathbf{y}(t)\ \cancel{\rightarrow}\ \mathbf{x}_{\text{{A}}}\big)\big|\mathbf{x}_{\text{{B}}}$ for each $t\in[0,T]$. ∎
@@
When $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{A}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})\equiv\mathbf{0}_{k_{\text{A}}\times k_{\text{B}}}$, the interpretation of “assigning infinite uncertainty to $\mathbf{x}_{\text{B}}$’s marginal likelihood” (Section A.4) becomes straightforward. In this case, $\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty$ simply corresponds to $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})^{-1}\to\mathbf{0}_{k_{\text{B}}\times k_{\text{B}}}$ for all $t$ and $\mathbf{x}$, since $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})$ becomes block-diagonal. For general turbulent systems or when this condition fails, the interpretation of this limiting procedure requires more care. Here, $\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty$ demands rigorous analysis of the structure of $\mathbf{x}_{\text{B}}$’s marginal likelihood and its impact on the Kalman gain operators. Only through such analysis can we properly nullify $\mathbf{x}_{\text{B}}$’s influence when testing for conditional assimilative causal links $\big(\mathbf{y}(t)\rightarrow\mathbf{x}_{\text{A}}\big)\big|\mathbf{x}_{\text{B}}$. We defer this detailed investigation to future work.
@@
When $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{A}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})\equiv\mathbf{0}_{k_{\text{A}}\times k_{\text{B}}}$ in a CGNS, implementing the condition within the generalized ACI framework $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})^{-1}\to\mathbf{0}_{k_{\text{B}}\times k_{\text{B}}}$ (Section A.4) leads to an equivalent reduced system. In this reduced system, the state estimation of $\mathbf{y}(t)$ becomes governed by the CGNS defined in (3.2a) and (3.2c), where $(\mathbf{x}_{\text{A}},\mathbf{y})$ form the state variables, with $\mathbf{x}_{\text{B}}$ being reduced to a deterministic forcing term defined by its observed values (analogous to a control or input term).
@@
This section includes three numerical studies that exploit the ACI framework to study complex dynamical systems with intermittency, non-Gaussian features, regime switching, and extreme events. All test cases employ CGNS models, leveraging their analytical tractability while capturing these nonlinear phenomena. Additionally, unless otherwise specified, the efficient approximation in (1.25) is used to compute the objective (C)CIR lengths.
D.1 A Nonlinear Dyad Model with Intermittent Extreme Events
@@
For completeness, we restate the nonlinear dyad model described in the main text (in differential form):
${\mathrm{d}}x$ $=(-d_{x}x+\gamma xy+f_{x}){\mathrm{d}}t+\sigma_{x}\mathrm{d}W_{x}$ (4.1a)
${\mathrm{d}}y$ $=(-d_{y}y-\gamma x^{2}+f_{y}){\mathrm{d}}t+\sigma_{y}\mathrm{d}W_{y}.$ (4.1b)
@@
This is a reduced-order conceptual model for atmospheric variability. It has been used to analyze the effects of various coarse-grained procedures on processes exhibiting intermittency, large-scale bifurcations, and microscale phase transitions. It is defined by an energy-conserving condition on its quadratic nonlinearities [majda2012physics]. The coupling parameter $\gamma>0$ plays a crucial role by ensuring significant positive $y$ values ($y>d_{x}/\gamma$) trigger extreme events in $x$.
@@
Supplementary Figure 1 presents the data assimilation results for the dyad model in (4.1). The key distinction between the filter and smoother distributions in estimating $y$ occurs prior to extreme events in $x$. This behavior is expected since the filter, operating without knowledge of future observations, cannot fully anticipate the triggering mechanism in $y$. In contrast, the smoother benefits from future extreme event data, enabling more accurate state estimation of y with reduced uncertainty.
@@
Supplementary Figure 1: Data assimilation of the dyad model (4.1). Panel (a): A single realization of the observed variable $x$. Panel (b): The true hidden signal $y$ (blue) alongside the posterior mean estimates from filtering (green) and smoothing (red), where the smoother is the complete smoother using all the information in future. The dashed line marks the anti-damping threshold, above which the net damping $d_{x}+\gamma y$ in (4.1a) becomes positive. Panel (c): Posterior variance of the filtered and smoothed estimates of $y$.
@@
Supplementary Figure 2 displays the ACI and CIR analyses. Several important patterns emerge:
@@
  • First, the ACI value (Panel (d)) reaches its maximum during $y$’s strongest anti-damping phase, corresponding to $y$’s peak instantaneous influence on $x$.
@@
  • Second, the subjective CIR attains its highest values (shown by the deep red shading in Panel (b)) slightly before $y>d_{x}/\gamma$ (initiating an extreme event) and for small $\varepsilon$ thresholds (indicating a longer-range influence). This reveals that extreme events develop gradually, with triggering conditions established well in advance. Notably, the objective CIR’s temporal extent does not reach the actual peak of extreme events. This finding mirrors the decorrelation time (the integration of the autocorrelation function) in complex dynamical systems, where the true causal influence often persists weakly beyond the formal timescale indicated by the objective function.
@@
  • Third, and most significantly, the objective CIR shortens as the system approaches $y$’s positive peak (corresponding to $x$’s extreme event buildup phase). At this stage, the filter can reliably detect the emerging pattern without requiring future information. This transition naturally partitions the time series at each extreme event, marking distinct dynamical regimes: a build-up phase with long-range dependence (where future information improves estimation) and an event phase where the trajectory becomes locally predictable. The short CIR also persists during $y$’s demise, where $x$ is the driving factor behind the system dynamics (high signal-to-noise ratio) and controls $y$ via $\gamma x^{2}$.
  • Finally, Panel (c) showcases the temporal evolution of the objective CIR for $y(t)\rightarrow x$, calculated using both the definition in (1.22) and its approximation in (1.25). This confirms the results of Theorem A.1 and (1.25), and clearly demonstrates the high skill that this computationally efficient approximation achieves while circumventing the expensive quadrature entailed by the $\varepsilon$-integrals of the subjective CIRs in the definition. Significantly, it captures the exact profile of the exact definition. This is true both during the quiescent phases of the signal, but importantly also at the onset of the extreme events generated by $y$.
@@
Supplementary Figure 2: ACI values and CIRs for the nonlinear dyad model (4.1) from $y$ to $x$ as functions of time. Panel (a): Time series of $x$ (magenta) and $y$ (blue), with the objective CIR depicted as whiskers extending forward in time from each $y(t)$. The dashed horizontal line marks the anti-damping threshold $d_{x}/\gamma$. Panel (b): Subjective CIR (shaded region, logarithmically scaled) as a function of the threshold $\varepsilon$ (logarithmic, reversed $y$-axis). Panel (c): Objective CIRs from $y$ to $x$ over time, calculated using both the definition (1.22) (in red) and the computationally efficient approximation (1.25) (in navy blue), corresponding to the whiskers in Panel (a). Panel (d): ACI values from $y$ to $x$ over time.
D.2 A Noisy Predator–Prey Model
The predator–prey model (also known as the Lotka-Volterra model) is fundamentally important across scientific disciplines as it captures the universal dynamics of interacting populations through simple yet powerful mathematics [cooke1981mathematical]. Originating in ecology to explain cyclical fluctuations between species like lynx and hares, its core principles have been adapted to model diverse systems, from disease spread in epidemiology to competition in economics and even chemical oscillations. The nonlinear feedback mechanisms in the model provide crucial insights into understanding stability, resilience, and emergent patterns of many natural and scientific problems.
Let us consider a noisy version of the predator–prey model:
${\mathrm{d}}x$ $=(\beta xy-\alpha x){\mathrm{d}}t+\sigma_{x}\mathrm{d}W_{x},$ (4.2a)
${\mathrm{d}}y$ $=(\gamma y-\delta xy){\mathrm{d}}t+\sigma_{y}\mathrm{d}W_{y},$ (4.2b)
@@ D.2 A Noisy Predator–Prey Model
Supplementary Figure 3 displays the data assimilation results. Panels (a)–(c) and (d)–(f) show the results by observing $y$ (recovering $x$) and observing $x$ (recovering $y$), respectively. The state estimation is more informative when the observed signal has a large value, corresponding to when the estimated variable induces an extreme event in the former’s evolution. Additionally, the uncertainty reduction in the smoother related to the filter is more significant when the predator $x$ is the observed variable, which accounts for the choice of the coupling, quadratic feedback parameter values; $\delta>\beta$.
Supplementary Figure 3: Data assimilation of the noisy predator–prey model (4.2). Panels (a)–(c) and (d)–(f) show the results by observing $y$ (recovering $x$) and observing $x$ (recovering $y$), respectively. The dashed lines in Panels (b) and (e) indicate the anti-damping threshold values.
Supplementary Figure 4 presents the ACI value (instantaneous causal strength) and CIR (influence duration) between predator ($x$) and prey ($y$) populations. The interaction exhibits two distinct regimes:
@@
Supplementary Figure 4: ACI values and CIRs for the noisy predator–prey model (4.2). Only the objective CIRs are shown. Panels (a)–(b) and (c)–(d) show the results from $x(t)$ to $y$ and from $y(t)$ to $x$, respectively. The dashed lines in Panels (a) and (b) indicate the anti-damping threshold values in the equations of $y$ and $x$, respectively.
D.3 A Stochastic Model Capturing the El Niño–Southern Oscillation (ENSO) Diversity
D.3.1 Model Details
@@
As was described in the main text, although few models can accurately capture ENSO diversity, a recently developed stochastic conceptual model successfully reproduces its diverse behaviors and non-Gaussian statistics [chen2022multiscale]. This model has been highlighted in a recent review [vialard2025nino], making it a suitable testbed for studying El Niño diversity. Using the defined abbreviations from the main text, the model consists of six state variables: ocean zonal current in the CP ($u$), WP thermocline depth ($h_{W}$), CP SST ($T_{C}$), EP SST ($T_{E}$), atmospheric winds ($\tau$, intraseasonal), and background Walker circulation ($I$, decadal). The variables ($u$, $h_{W}$, $T_{C}$, and $T_{E}$) operate on interannual timescales and model the anomalies away from the corresponding climatology. As a nonlinear system with state-dependent noise, the model generates extreme events and intermittency. Note that the decadal variable represents the strength of the Walker circulation. The two SST variables ($T_{C}$ and $T_{E}$) allow reconstruction of spatiotemporal SST patterns across the equatorial Pacific, providing an intuitive way to identify different ENSO event types. The model reads:
${\mathrm{d}}u$ $=\big(-ru-\delta_{u}\frac{T_{C}+T_{E}}{2}+\beta_{u}(I)\tau\big)\mathrm{d}t+\sigma_{u}\mathrm{d}W_{u},$ (4.3a)
${\mathrm{d}}h_{W}$ $=\big(-rh_{W}-\delta_{h}\frac{T_{C}+T_{E}}{2}+\beta_{h}(I)\tau\big)\mathrm{d}t+\sigma_{h}\mathrm{d}W_{h},$ (4.3b)
${\mathrm{d}}T_{C}$ $=\big(\left(r_{C}-c_{1}(t,T_{C})\right)T_{C}+\zeta_{C}T_{E}+\gamma_{C}h_{W}+\sigma(I)u+C_{u}+\beta_{C}(I)\tau\big)\mathrm{d}t+\sigma_{C}\mathrm{d}W_{C},$ (4.3c)
${\mathrm{d}}T_{E}$ $=\big(\left(r_{E}-c_{2}(t)\right)T_{E}-\zeta_{E}T_{C}+\gamma_{E}h_{W}+\beta_{E}(I)\tau\big)\mathrm{d}t+\sigma_{E}\mathrm{d}W_{E},$ (4.3d)
${\mathrm{d}}\tau$ $=-d_{\tau}\tau\mathrm{d}t+\sigma_{\tau}(t,T_{C})\mathrm{d}W_{\tau},$ (4.3e)
${\mathrm{d}}I$ $=-\lambda(I-m)\mathrm{d}t+\sigma_{I}(I)\mathrm{d}W_{I}.$ (4.3f)
@@
While the full model details are available in [chen2022multiscale], its basic mechanism can be summarized as follows. EP El Niño events are typically associated with a strong thermocline buildup in the WP, whereas CP EN and La Niña events are primarily triggered by advective processes. In the latter case, when $\sigma(I)\propto I$ is larger, indicating a strengthening Walker circulation, the advection term $u$ becomes dynamically significant in driving $T_{C}$. Unlike the standard discharge–recharge oscillator model, which couples $h_{W}$ and $T_{E}$, the $T_{C}$ variable acts as a transitional component that is expected to have a more direct influence on $T_{E}$.
D.3.2 Additional Results Using Model-Generated Data